If the energy of electron in second excited state is 9eV in an infinite potential box . The ground state energy of the particle is:


1
Expert's answer
2021-12-29T12:34:55-0500

"E_2=E_0(1-\frac 1{2^2}),\implies"

"E_0=\frac 43 E_2=\frac 43\cdot 9=12~eV."